What precipitate is most likely formed from a solution containing Ba+2, Na+1, OH-1, and CO3-2.
1
Expert's answer
2019-06-05T03:58:34-0400

Barium carbonate salt will form as a slightly soluble precipitate:

Ba+2(aq) + 2OH-1(aq) + 2Na+1(aq) + CO3-2(aq) = BaCO3(s) + 2Na+1(aq) + 2OH-1 (aq)


The equation above is an ionic equation of the following chemical reaction:

Ba(OH)2 (aq) + Na2CO3 (aq)= BaCO3 (s) +2NaOH (aq)


where formation of a precipitate is a driving force of the reaction.